OpenStudy (anonymous):

-3(2x-6)+x=5(x+2) solve for x

OpenStudy (anonymous):

x = 4/5

OpenStudy (anonymous):

-3(2x-6)+x=5(x+2 -(3 * 2x - 3 * 6) + x = 5(x + 2) -(6x - 18) + x = 5(x + 2)